Find two consecutive even numbers whose sum is 22.
Found 2 solutions by algebrahouse.com, Alan3354:
Answer by algebrahouse.com(1659)

Find two consecutive even numbers whose sum is 22.

x = 1st consecutive even number
x + 2 = 2nd consecutive even number

x + x + 2 = 22 {their sum is 22}
2x + 2 = 22 {combined like terms}
2x = 20 {subtracted 2 from both sides}
x = 10 {divided both sides by 2}
x + 2 = 12 {substituted 10, in for x, into x + 2}

10 and 12 are the two consecutive even numbers

For more help from me, visit: www.algebrahouse.com



Answer by Alan3354(69443)

22/2 = 11, the center or average
--> 10 & 12